### v2

# Replayable Financial Agents: A Determinism-Faithfulness Assurance Harness for Tool-Using LLM Agents

Abstract
LLM agents struggle with regulatory audit replay: when asked to reproduce a flagged transaction decision with identical inputs, most deployments fail to return consistent results. We introduce the Determinism-Faithfulness Assurance Harness (DFAH), a framework for measuring trajectory determinism and evidence-conditioned faithfulness in tool-using agents deployed in financial services.
[…]
Code and data: https://github.com/ibm-client-engineering/output-drift-financial-llms
1 Introduction
LLM agents are being adopted rapidly for financial decision-making tasks including compliance triage, portfolio rebalancing, and regulatory reporting (Xing, 2025). For many institutions, performance of these agents plays a role in allowing them to scale operations, reduce manual review burden, or accelerate time-sensitive workflows. Therefore, attainment of a more complete understanding of the reproducibility characteristics of these agents would be pertinent to practitioners navigating regulatory requirements.
When a regulator asks “Why did your system flag this transaction?” or “What evidence supported this investment decision?”, institutions must demonstrate two things: (1) that the decision can be replayed with identical inputs to produce the same output (determinism), and (2) that the decision was tied to retrieved evidence rather than fabricated reasoning (faithfulness). Our recent work has quantified output drift in LLM labeling tasks (Khatchadourian and Franco, 2025), but the agentic setting introduces additional complexity given that agents make multi-step tool calls, and drift can occur at the trajectory level (different tool sequences) even when final decisions match.
This is not to say that a consensus has been reached on how to measure these properties however, as Anthropic’s recent engineering guidance notes that “the capabilities that make agents useful—autonomy, intelligence, flexibility—also make them harder to evaluate” (Anthropic, 2026). Financial services has responded with “skills”—procedural wrappers that promise deterministic execution—but the question of whether deployed agents actually achieve this determinism remains open. The DFAH framework provides evaluation infrastructure to answer it.
In this study, we attempt to formalize trajectory determinism, decision determinism, and evidence-conditioned faithfulness for tool-using agents while simultaneously examining the relationship between these metrics across model sizes and architectures. Given the fact that this topic has significance to practitioners in regulated environments, there exists a growing body of work applying evaluation techniques to find similar relationships (Ludwig et al., 2024; Halperin, 2025). We feel our results will provide further validation to emerging patterns and seek to expand on existing research by incorporating stress-test scenarios (deployment cycles, stale data, data-quality faults, market shocks) and three financial benchmarks with ground-truth labels. Additionally, our framework aligns with the trial/trajectory/grader terminology now standard in production deployments, including the pass@k@k vs passk distinction relevant to compliance contexts.
Contributions.
  1. 1.
    DFAH framework: Formal definitions of trajectory determinism, decision determinism, and evidence-conditioned faithfulness for tool-using agents, with open-source implementation.
  2. 2.
    Empirical findings: Determinism and task accuracy show no detectable correlation (r=0.11r=-0.11, 95% CI [0.49,0.31][-0.49,0.31], $n=21$), indicating that single-metric evaluation is insufficient—both dimensions must be measured independently.
  3. 3.
    Large-scale agentic evaluation: 4,700+ runs across 7 models (4 providers), 3 financial benchmarks (compliance triage, portfolio constraints, DataOps exceptions; 50 cases each), revealing distinct determinism profiles across model tiers.
  4. 4.
    Determinism-accuracy tradeoff characterization: Small models achieve near-perfect determinism through pattern matching (low accuracy), while frontier models reason through diverse tool paths (higher accuracy, lower determinism).
  5. 5.
    Practical guidance: Model tier recommendations, validation scaling factors, and stress-test protocols for compliance-critical deployments.
2 Related Work
Prior research has consistently linked LLM output variability with model architecture and task structure, though often without consistent evaluation methodology. The ML reproducibility literature (Pineau et al., 2021; Kapoor and Narayanan, 2023) has established that evaluation choices—data splits, random seeds, hyperparameters—can substantially alter reported results. In prior work (Khatchadourian and Franco, 2025), we established that LLM outputs vary across runs even at temperature zero, with drift rates depending on model size and task characteristics. Ouyang et al. (2024) document similar non-determinism in code generation, finding that identical prompts produce varying outputs even at temperature zero. The agentic setting introduces additional complexity, as agents make multi-step tool calls and drift can compound at the trajectory level even when final decisions match.
Research such as that conducted by Ludwig et al. (2024) provides a framework for using LLM-generated labels in econometric research, showing how validation subsamples can debias coefficient estimates. We adopt their validation scaling approach and extend it with drift-adjusted sample sizing. Halperin (2025) introduce information-theoretic metrics for measuring semantic faithfulness, using Kullback-Leibler divergence to quantify how well answers align with provided context. Their entropy production measures provide principled approaches to detecting structurally unfaithful outputs, motivating the evidence-conditioned faithfulness metric in this work.
In the financial LLM space, Wu et al. (2023) demonstrate domain-specific pre-training on financial data, while Yang et al. (2023) provide open-source alternatives. Both focus on task performance rather than reproducibility—our work is complementary, measuring whether these models produce consistent outputs when deployed as agents in regulated workflows.
Wang and Wang (2025) provide the first comprehensive assessment of LLM output consistency in finance and accounting research, finding that binary classification achieves near-perfect reproducibility while complex tasks exhibit greater variability. Their work establishes that downstream statistical inferences remain robust despite measurable output variance—a finding adequate for research validity but insufficient for regulatory examination where instance-level reproducibility is required. The DFAH framework extends this analysis to tool-using agents, where trajectory-level variance introduces additional complexity beyond single-output consistency.
The consistency of these relationships across evaluation contexts is less transparent however, as OpenAI’s work on chain-of-thought monitoring (OpenAI, 2025) demonstrates that reasoning traces can be misleading or unfaithful to actual model computations. This reinforces the need for evidence-alignment checks that verify decisions against retrieved context rather than generated reasoning—a deterministic agent with opaque internal reasoning is preferable to a non-deterministic agent producing plausible-sounding but unverifiable explanations.
Several benchmarks evaluate LLM agents on multi-step tasks: SWE-bench (Jimenez et al., 2024) measures code-level problem solving, AgentBench (Liu et al., 2024) evaluates across diverse environments, and GAIA (Mialon et al., 2024) tests general assistant capabilities. These focus on task completion; our work instead measures consistency of that completion across repeated runs—a requirement absent from these benchmarks but central to regulated deployments. Tool-augmented language models (Schick et al., 2023; Yao et al., 2023; Patil et al., 2023) have shown that external tool use can improve factual grounding, but also introduce trajectory-level variance not present in single-output generation.
Anthropic’s engineering guidance on agent evaluation (Anthropic, 2026) establishes terminology now becoming industry standard: a trial is a single attempt at a task, a transcript is the complete record of that attempt, and a grader is logic scoring performance. The DFAH framework implements this structure with financial-domain graders for determinism and faithfulness. Emerging work suggests that smaller, task-optimized models can match larger systems on specific benchmarks—Gan et al. (2025) show that Qwen2.5-7B fine-tuned on motion-specific data matches or exceeds larger models on video motion understanding. The observation that 7–20B models achieved higher determinism than frontier models on these financial benchmarks is consistent with this pattern, although domain-specific validation remains necessary.
Song et al. (2025) evaluate LLMs across the scientific discovery pipeline, finding performance gaps on discovery-focused tasks and noting systematic weaknesses across scientific domains. This observation is relevant for financial applications where model limitations may manifest as plausible-sounding but poorly-supported recommendations. Saini et al. (2025) address AI-driven data quality control in regulated financial environments, emphasizing audit trails and configurable policies—the DataOps Exception benchmark directly addresses this use case.
3 Problem Formulation
In order to formalize what “replayable” means in the context of regulatory audit, we first establish definitions connecting audit requirements to measurable properties.
3.1 Agent Model
Consider a tool-using LLM agent $\mathcal{A}$ that, given an input query $q$ and access to tools $\mathcal{T}=\{t_{1},\ldots,t_{k}\}$, produces:
  • A trajectory $\tau=[(t_{i_{1}},a_{1},r_{1}),\ldots,(t_{i_{n}},a_{n},r_{n})]$ of tool calls with arguments $a_{j}$ and results $r_{j}$;
  • A decision $d\in\mathcal{D}$ where $\mathcal{D}$ is task-specific (e.g., {escalate, dismiss, investigate} for compliance triage).
Following the terminology from Anthropic (2026), each execution constitutes a trial, the trajectory $\tau$ is the transcript, and the determinism/faithfulness checks are graders.
3.2 Determinism Metrics
Given $N$ independent runs of $\mathcal{A}$ on input $q$, producing trajectories $\{\tau^{(1)},\ldots,\tau^{(N)}\}$ and decisions $\{d^{(1)},\ldots,d^{(N)}\}$, comparisons are anchored to run (1)—matching audit replay where a recorded trace serves as the reference execution:
Definition 1 (Action Determinism).
The fraction of runs with identical tool sequences:
$\text{ActDet}(q)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}[\text{tools}(\tau^{(i)})=\text{tools}(\tau^{(1)})]$
Definition 2 (Signature Determinism).
The fraction of runs with identical tool sequences and arguments:
$\text{SigDet}(q)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}[\tau^{(i)}=\tau^{(1)}]$
Definition 3 (Decision Determinism).
The fraction of runs with identical final decisions:
$\text{DecDet}(q)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}[d^{(i)}=d^{(1)}]$
Run-Level vs. Case-Level Aggregation.
For audit replay validation, two aggregation methods are used:
  • Run-Level Determinism: Average fraction of individual runs matching the reference (Definition 3 above). For $n$ test cases with $k$ runs each, this aggregates over all $n\times k$ runs. Yields continuous values.
  • Case-Level Determinism: Fraction of test cases where all $k$ runs produce identical results. For $n$ cases, valid values are multiples of $1/n$ (e.g., 0%, 10%, 20%, …, 100% for $n=10$). This stricter metric reflects the passk audit requirement.
Benchmark tables report case-level determinism (the stricter audit metric), while stress tests report run-level determinism for finer-grained comparison.
For audit purposes, decision determinism is primary: an auditor cares that the agent reaches the same conclusion, even if internal reasoning paths vary.
3.3 Pass@k vs Passk: Why Compliance Needs the Stricter Metric
Agent evaluation literature distinguishes two success metrics with opposite scaling behaviors (Anthropic, 2026), and the choice between them has significant implications for compliance validation:
Definition 4 (Pass@k (Optimistic)).
The probability of achieving at least one success in $k$ trials:
$\text{pass}@k=P(\text{at least one success in }k\text{ trials})$
As $k$ increases, pass@k@k approaches 100%—more attempts mean higher odds of finding one working solution.
Definition 5 (Passk (Conservative)).
The probability that all $k$ trials succeed:
$\text{pass}^{k}=P(\text{all }k\text{ trials succeed})$
As $k$ increases, passk falls—demanding consistency across more trials.
Proposition 1.
For compliance-critical deployments, passk is the relevant metric.
Rationale: A regulatory examiner selecting $k$ random historical decisions for replay expects all $k$ to reproduce correctly. The pass@k@k metric—where one success suffices—applies to software engineering contexts (e.g., code generation where any working solution is acceptable). Financial compliance requires passk: every decision must be reproducible, not just some of them.
The framework directly measures passk through trajectory determinism. Tier 1 models achieving 100% determinism have passk = pass@1@1 for all $k$: every trial produces identical results, so the metric is invariant to $k$. This is the ideal compliance posture.
3.4 Faithfulness Metrics
Determinism is necessary but not sufficient—an agent that consistently produces the same wrong answer is still wrong. Therefore we also measure faithfulness, defined as the degree to which decisions are tied to retrieved evidence rather than fabricated reasoning.
Following Halperin (2025), faithfulness is not a measure of “truth” (which would require ground truth labels). Rather, it measures alignment relative to retrieved artifacts, which is auditable:
Definition 6 (Evidence Grounding).
Let $E=\{e_{1},\ldots,e_{m}\}$ be the evidence retrieved during trajectory $\tau$, and let $C=\{c_{1},\ldots,c_{p}\}$ be claims made in the decision rationale. Then:
$\text{EvidGround}(\tau,d)=\frac{|\{c_{i}:\exists e_{j}\text{ s.t. }c_{i}\sqsubseteq e_{j}\}|}{|C|}$
where $c_{i}\sqsubseteq e_{j}$ denotes that claim $c_{i}$ is aligned with evidence $e_{j}$ (defined below).
Evidence-Alignment Heuristic.
The alignment relation $c_{i}\sqsubseteq e_{j}$ is operationalized as a lexical/semantic heuristic (not NLI entailment): (1) extract noun phrases and numerical values from both claim and evidence, (2) compute token-level Jaccard similarity, (3) flag as aligned if similarity $\geq 0.6$ AND key entities match. For numerical claims (amounts, dates, thresholds), exact match within tolerance is required.
This heuristic prioritizes interpretability over statistical power. Embedding-based, NLI, or LLM-as-a-judge approaches introduce recursive non-determinism—using an LLM to evaluate faithfulness defeats the purpose of an audit harness. False negatives (undercounting aligned claims due to paraphrase) are acceptable; false positives (masking fabricated claims) are the regulatory failure mode.
This conservative approach ensures that any claim marked as “aligned” can be traced to specific evidence by human reviewers using inspectable rules. Validation on a 100-sample slice from Compliance Triage achieved 89% agreement with human annotation.
Calibration and Limitations.
The heuristic is designed for high precision, low recall: it reliably identifies claims with clear lexical grounding but may undercount valid paraphrased alignments. This design choice means faithfulness scores reported here are likely conservative—true evidence grounding may be higher than measured. The 89% human agreement was validated on Compliance Triage; other tasks with different linguistic patterns (e.g., Portfolio Constraint with numerical reasoning) may show different calibration. Correlations involving faithfulness should be interpreted with this measurement bias in mind, particularly for schema-templated outputs where lexical overlap is structurally high.
Definition 7 (Constraint Satisfaction).
For decisions subject to constraints $\mathcal{K}=\{k_{1},\ldots,k_{r}\}$ (e.g., position limits, regulatory thresholds):
$\text{ConSat}(d)=\frac{|\{k\in\mathcal{K}:d\text{ satisfies }k\}|}{|\mathcal{K}|}$
4 Evaluation Framework
The Determinism-Faithfulness Assurance Harness implements the evaluation infrastructure described by Anthropic (2026), adapted for financial compliance requirements.
4.1 Harness Architecture
The DFAH comprises four components:
Task Runner:
Executes agent trials with controlled randomness (T=0.0, seed=42 where supported), recording full transcripts.
Trajectory Store:
Persists all tool calls, arguments, and results in structured format for replay and comparison.
Grader Suite:
Three grader types following Anthropic (2026):
  • Code-based: Tool call verification, schema validation, decision matching (deterministic, fast)
  • Model-based: Faithfulness assessment via LLM judge (flexible but introduces variance)
  • Human: Audit sampling for calibration (gold standard, expensive)
Aggregator:
Computes tier-level statistics, validation scaling factors, and deployment recommendations.
For production use, code-based graders are recommended for determinism checks (they introduce no additional variance) and human graders for faithfulness calibration. In the reported experiments, determinism was evaluated with code-based graders and faithfulness with the heuristic evidence-grounding metric (Section 3.4) plus human calibration. Model-based graders are included in the architecture for completeness but were not used in the study’s reported results; they are suitable for development iteration where speed matters more than precision.
4.2 Stress-Test Scenarios
The harness evaluates agents under four stress scenarios designed to simulate real-world disruptions:
Redeploy Perturbation:
Same prompt, fresh agent instance after container restart or model reload. Tests intra-model consistency across deployment cycles—critical for production environments where models are regularly redeployed.
Data Quality Fault Injection:
Inject {NULL, NaN, outlier} values into 10% of tool responses. Simulates real-world data pipeline failures and tests error handling robustness.
Temporal Shift:
Use context data from 6 months prior (stale filings, outdated market data). Tests whether agents correctly identify and handle outdated information. [^2] Temporal Shift perturbation not included in the experiments reported here; planned for future work.
Market Shock Simulation:
Modify numeric tool outputs by $\pm 3\sigma$ to simulate extreme volatility. Tests decision stability under stress conditions.
For each perturbation $p$ and agent configuration $(m,a)$:
$\Delta\text{Det}_{p}(m,a)=\text{DecDet}_{\text{baseline}}-\text{DecDet}_{p}$
A robust agent configuration should maintain $\Delta\text{Det}_{p}<10\%$ across all perturbation types.
Projection Methodology.
For configurations where full stress-scenario runs were not completed, stress determinism is projected as:
$\widehat{\text{Det}}_{p}(m,a)=\text{Det}_{\text{baseline}}(m,a)-\Delta_{\text{tier},p}$
where $\Delta_{\text{tier},p}$ is the median degradation observed for that model tier under perturbation $p$ across pilot configurations (e.g., $\Delta_{\text{Tier1,DQ}}\approx 6\%$, $\Delta_{\text{Frontier,Redeploy}}\approx 12\%$). In Table 9, baseline values are measured; stress columns marked are projected using this formula.
4.3 Architecture Variants
Two agent architectures are evaluated:
Unconstrained:
Standard ReAct-style agent with free-form reasoning
Schema-First:
Structured output with JSON schemas for all decisions
The schema-first architecture aligns with the “skills” pattern emerging in production deployments, where deterministic code execution wraps LLM capabilities. The experiments validate that this architectural pattern measurably improves determinism without sacrificing task performance.
4.4 Validation Sample Sizing
Following Ludwig et al. (2024), the required validation sample size to achieve target precision $\epsilon$ in debiased estimates:
$n_{\text{val}}^{*}=\left(\frac{\sigma_{\text{drift}}}{\epsilon}\right)^{2}\cdot\phi(m)$
where $\phi(m)$ is a model-specific scaling factor derived from drift variance. Tier 3 models require $\phi(m)\approx 3.7\times$ the validation samples of Tier 1 models to achieve equivalent statistical reliability.
5 Benchmark Tasks
Three benchmark tasks representative of financial agent use cases are provided (Table 1). Each task includes 50 test cases with ground-truth labels, tool definitions, and evaluation scripts.
Table 1: Benchmark tasks for financial LLM agent evaluation. Each task includes 50 test cases with ground-truth labels and expected tool sequences.
Task Domain Decisions Tools
Compliance Triage Regulatory escalate/dismiss/investigate 3
Portfolio Constraint Investment approve/reject/modify 5
DataOps Exception Operations auto_fix/escalate/quarantine 6
5.1 Compliance Triage Benchmark
The compliance triage task evaluates an agent’s ability to process transaction alerts and decide whether to escalate, dismiss, or investigate.
Tool Set.
  • check_sanctions_list(entity_name) $\rightarrow$ {match_status, match_score, list_type}
  • get_customer_profile(customer_id) $\rightarrow$ {risk_rating, kyc_status, account_age, …}
  • calculate_risk_score(transaction_id) $\rightarrow$ {score, factors, threshold_status}
Decision Space.
$\mathcal{D}=\{\text{ESCALATE},\text{DISMISS},\text{INVESTIGATE}\}$
Alert Categories.
The 50 test alerts span:
  • Sanctions-related (8): Direct hits, sanctions-adjacent entities, high-risk destinations
  • PEP exposure (4): Politically exposed persons, FCPA risk
  • Structuring patterns (6): Just-under-threshold transactions, velocity spikes
  • High-value goods (7): Art, jewelry, luxury vehicles, real estate
  • High-risk sectors (10): Gaming, weapons, crypto, extractive industries
  • Standard business (15): Intercompany transfers, vendor payments (dismiss baseline)
Ground Truth.
15 escalate, 25 dismiss, 10 investigate.
Determinism Requirements.
For an agent to pass compliance audit:
  • Decision determinism $\geq 95\%$ for escalate cases (high-risk alerts)
  • Trajectory determinism $\geq 90\%$ for all cases (audit replay)
  • Evidence grounding $\geq 80\%$ for all decisions
5.2 Portfolio Constraint Benchmark
The portfolio constraint task validates proposed trades against position limits, sector caps, and liquidity requirements.
Tool Set.
  • get_portfolio_positions(portfolio_id) $\rightarrow$ {holdings, weights, sectors}
  • get_market_data(ticker) $\rightarrow$ {price, volume, volatility, beta}
  • check_concentration_limits(portfolio_id, ticker, proposed_weight) $\rightarrow$ {compliant, limit, current, proposed}
  • calculate_var(portfolio_id, horizon, confidence) $\rightarrow$ {var, cvar, contribution}
  • get_regulatory_constraints(account_type) $\rightarrow$ {constraints, thresholds}
Constraint Types.
  • Position limit violations (12 cases): Exceeding 5% single-stock concentration
  • Sector cap violations (6 cases): Technology sector exceeding 25% cap
  • Liquidity violations (8 cases): Small-cap stocks with insufficient volume
  • Cash reserve violations (3 cases): Depleting below 2% minimum
  • Clean approvals (21 cases): Trades satisfying all constraints
Ground Truth.
25 approve, 18 reject, 7 modify.
5.3 DataOps Exception Benchmark
The DataOps task evaluates exception resolution in financial data pipelines, addressing AI-driven QC requirements in regulated data operations (Saini et al., 2025).
Tool Set.
  • get_exception_details(exception_id) $\rightarrow$ {type, field, value, rule_violated}
  • query_reference_data(field, value) $\rightarrow$ {valid, canonical_value, alternatives}
  • get_historical_fixes(exception_type) $\rightarrow$ {fixes, success_rates, avg_resolution_time}
  • validate_fix(exception_id, proposed_fix) $\rightarrow$ {valid, conflicts, warnings}
  • apply_fix(exception_id, fix) $\rightarrow$ {success, audit_trail}
  • escalate_to_human(exception_id, reason) $\rightarrow$ {ticket_id, assigned_to}
Exception Types.
  • Format errors (15): Date formats, currency codes, numeric parsing
  • Business rule violations (18): Negative prices, invalid ratios
  • Reference data mismatches (10): Ticker mappings, CUSIP/ISIN lookups
  • Missing required fields (7): CUSIPs, settlement dates
Ground Truth.
30 auto_fix, 15 escalate, 5 quarantine.
6 Experiments
6.1 Dataset
The evaluation uses 74 configurations across 12 models from 4 providers. Models are classified into tiers based on observed determinism behavior at T=0.0, not solely parameter count:
  • Tier 1 (target 100% det.): qwen2.5:7b, gpt-oss:20b, deepseek-r1:8b, granite-3-8b-ollama (36 configs, all 7–20B params, local inference)
  • Tier 2 (50–90% det.): llama-3-3-70b, granite-3-8b-watsonx (20 configs, cloud inference with potential API-level variance)
  • Tier 3 (<<20% det.): gpt-oss-120b (6 configs, 120B params)
  • Frontier: claude-opus-4.5, gemini-2.5-pro (12 configs, API) [^3] Model identifiers throughout use marketing names (e.g., claude-opus-4.5). For reproducibility, the corresponding API model IDs are: claude-opus-4-5-20250514, claude-sonnet-4-20250514, gemini-2.0-flash, gemini-2.5-pro. Local Ollama models were pulled as qwen2.5:7b-instruct, granite3.3:8b, deepseek-r1:8b, gpt-oss:20b.
Note: granite-3-8b appears in both Tier 1 (local Ollama) and Tier 2 (watsonx cloud) because cloud API inference introduces additional variance. Tiering is based on non-agentic experiments; agentic tool-use can introduce additional system variance.
Statistical Precision.
At $n=8$ runs per configuration and $\alpha=0.05$, effect sizes $\geq$40% are reliably detectable (e.g., distinguishing 100% from 60% determinism). The design prioritizes detecting instability (any non-zero drift) rather than precise effect estimation. For binomial proportions with $n=8$:
  • 100% observed (8/8): 95% CI [63%, 100%] (Wilson)
  • 90% observed (7/8): 95% CI [56%, 100%]
  • Tier aggregates use $n=24$ runs for tighter intervals
Consequently, 100% observed determinism (8/8 identical) confirms no drift was detected at this run count; rare drift occurring <<5% of runs may go unobserved.
6.2 Results
Data provenance: Tables 46 report measured results from v2 agentic experiments (4,705 runs across 7 models, March 2026). Table 2 retains v1 non-agentic baseline data (74 configurations, December 2025) for continuity with prior results. Table 9 reports v1 pilot stress-test data (10-case subset) with projected values marked .
Model Tier Analysis.
Table 2 summarizes determinism and validation scaling by model tier.
Table 2: Model tier analysis across 74 configurations (non-agentic baseline at T=0.0T{=}0.0). Tiers are defined by decision determinism in structured tasks without tool use. Agentic benchmarks with tool calling may show different rates (see Table 4). Validation scaling indicates the factor by which validation samples must increase relative to Tier 1 baseline to achieve equivalent statistical reliability.
[…]
Figure 1: Decision determinism by model across agentic benchmarks at T=0.0T{=}0.0 (averaged over all benchmarks per model). Tier 1 local models (7–8B) achieve 91–98% determinism via pattern matching but low accuracy; the 20B GPT-OSS model drops to 77%; frontier API models range from 59% (Gemini 2.5 Pro) to 86% (Gemini 2.0 Flash). The red dashed line marks the 95% compliance threshold—only Qwen 2.5 exceeds it, but at the cost of 33% accuracy.
Determinism-Accuracy Analysis.
Terminology note: The framework defines faithfulness as evidence grounding (Section 3.4), a measure of whether decisions cite retrieved evidence. Tables 46 report accuracy—majority-vote agreement with ground-truth labels—which is a distinct construct. This section analyzes the determinism-accuracy relationship; faithfulness is not evaluated in the v2 agentic experiments due to the cost of human annotation at scale.
Correlation analysis is conducted over $n=21$ model-benchmark configurations derived from the measured v2 agentic benchmark suite; this is distinct from the 74-configuration v1 non-agentic baseline retained in Table 2 for historical comparison.
Preliminary analysis in v1 suggested a positive correlation ($r=0.45$) between determinism and faithfulness based on reconstructed data. Expanded v2 experiments with measured data across $n=21$ model-benchmark configurations found no significant relationship: Pearson r=0.11r=-0.11, 95% BCa CI [0.49,0.31][-0.49,0.31], $p=0.63$ (10,000 bootstrap resamples, seed=42; Efron and Tibshirani, 1994). Spearman rank correlation confirmed the null result ($\rho_{s}=-0.08$, 95% CI [0.50,0.38][-0.50,0.38], $p=0.74$). Partial correlation controlling for $\log(\text{model size})$, task type, and inference environment yielded $r_{partial}=-0.01$, 95% CI [0.49,0.42][-0.49,0.42], $p=0.98$.
The v1-to-v2 discrepancy reflects two methodological improvements: (1) v2 uses standardized evaluation conditions (T=0.0T{=}0.0, structured system prompt) across all models, whereas v1 conflated trajectory and decision determinism; (2) v2 replaces reconstructed data with direct measurements from 4,705 agentic runs, revealing that the apparent v1 correlation was an artifact of simulated tier-level aggregation.
This null correlation supports the case for DFAH. If determinism and accuracy were correlated, measuring one could serve as a proxy for the other. Because neither predicts the other in our data, both must be measured separately—precisely what DFAH provides. The data reveal three distinct operational profiles:
  • High determinism, low accuracy: qwen2.5:7b achieves 94–100% decision determinism but 20–42% accuracy by rigidly pattern-matching to default actions (e.g., “investigate” 76% of compliance cases).
  • Moderate determinism, higher accuracy: Claude Sonnet achieves 82–86% decision determinism with 14–67% accuracy, exploring diverse tool paths (26–58% signature determinism).
  • Variable determinism, variable accuracy: Gemini 2.5 Pro shows 50–68% decision determinism with 40–57% accuracy, demonstrating the widest behavioral variance.
No model occupies the “high determinism and high accuracy” quadrant, underscoring that determinism alone is insufficient for deployment—a finding that directly motivates DFAH’s multi-dimensional measurement approach.
Figure 2: Decision determinism vs. task accuracy across $n=21$ model-benchmark configurations at T=0.0T{=}0.0 (r=0.11r=-0.11, 95% BCa CI [0.49,0.31][-0.49,0.31], $p=0.63$). The null correlation indicates no detectable relationship between determinism and accuracy in our sample: small models (upper-left) achieve high determinism via pattern matching, frontier models (center) show moderate determinism with variable accuracy, and no model achieves both perfect determinism and high accuracy. Two partial-coverage models (deepseek-r1:8b, mistral:7b; 2 compliance cases each) are included as data points but omitted from Tables 46 due to insufficient coverage.
Task-Structure Effect.
Determinism varies significantly by task structure (Table 3):
Table 3: Task-structure effect on determinism at T=0.0 (non-agentic v1 baseline tasks).
Task Det. Mean (%) Det. Std n
SQL Generation 92.7 23.6 20
Summarization 90.1 26.6 17
RAG Retrieval 75.9 29.6 14
Figure 3: Task-structure effect on determinism at T=0.0T{=}0.0 across all models. Decision determinism (solid bars) and signature determinism (translucent bars) shown per benchmark. Error bars indicate standard deviation across models. Compliance Triage (82.3% decision, 69.6% signature) and DataOps Exception (84.0%, 56.8%) show similar decision determinism but divergent signature determinism. Portfolio Constraint (79.2%, 64.0%) shows the lowest decision determinism due to multi-step constrained reasoning. No benchmark exceeds the 95% compliance threshold on average.
The agentic benchmarks (Tables 46) show a related but distinct pattern: decision determinism is comparable across financial tasks (79–84%), but signature determinism—requiring identical tool-call sequences—drops substantially for semi-structured tasks (DataOps: 56.8%), confirming that tool-path variance—not decision variance—is the primary source of non-reproducibility in agentic settings.
Agentic Benchmark Results.
Table 4 shows determinism from agentic experiments on Compliance Triage (50 alerts, 3–8 runs per model per case). Agentic tool use introduces additional system variance beyond the non-agentic baselines in Table 2. Seven models spanning local inference (Ollama) and API providers (Anthropic, Google) are evaluated.
Table 4: Agentic compliance triage benchmark results (50 cases at T=0.0T{=}0.0). Dec.Det = case-level decision determinism (fraction of 50 alerts where all runs produced identical decisions); Sig.Det = signature (tool sequence) determinism (case-level); Acc = majority-vote accuracy against ground-truth labels. For run-level architecture comparisons, see Table 9.
Model Tier Dec.Det (%) Sig.Det (%) Acc (%) Tools/Run
qwen2.5:7b Tier 1 100.0 100.0 38.0 2.9
gpt-oss:20b Tier 1 86.0 92.0 34.8 4.0
gemini-2.0-flash Frontier 68.0 66.0 52.0 3.8
gemini-2.5-pro Frontier 68.2 40.9 40.2 3.3
claude-sonnet-4 Frontier 82.0 44.0 33.3 4.6
claude-opus-4.5 Frontier 72.0 44.0 68.7 4.2
Text-only baseline (no tool invocation):
granite3.3§ Tier 1 100.0 100.0 34.0 0.0
§granite3.3 does not invoke tools; decisions are text-only (not comparable to tool-calling models).
Note: The benchmarks exist to generate decisions that can be measured for reproducibility, not to test whether models are good at compliance triage. Accuracy varies by model capability; the focus is determinism (reproducibility for audit).
Tier 1 models (qwen2.5:7b, granite3.3) achieve perfect decision determinism but through pattern matching—qwen defaults to “investigate” on 76% of cases. Frontier models show a different profile: claude-opus-4.5 achieves the highest accuracy (68.7%) with 72% decision determinism, while its tool paths vary substantially (44% signature determinism).
Observation: Same Conclusion, Different Reasoning.
Across all three benchmarks, frontier models exhibit a consistent pattern: moderate-to-high decision determinism (71–86%) paired with substantially lower signature determinism (20–58%). This gap indicates that these models converge on the same final decisions while exploring meaningfully different tool-call sequences across runs. For audit purposes, this pattern is consequential—it means decision replay succeeds while trajectory replay fails, requiring evaluators to distinguish which level of reproducibility their regulatory context demands.
Decision Concentration Bias.
Small models exhibit extreme concentration in their decision distributions: qwen2.5:7b selects “investigate” for 76% of compliance alerts and “modify” for 82% of portfolio trades, regardless of case content. This yields high determinism as an artifact of low decision entropy rather than consistent reasoning. We quantify this via the normalized entropy of the decision distribution $H(\mathbf{d})/\log|\mathcal{D}|$, where values near 0 indicate concentration on a single action and values near 1 indicate uniform distribution across the decision space $\mathcal{D}$. For compliance triage ($|\mathcal{D}|=3$): qwen2.5:7b has $H/\log|\mathcal{D}|=0.56$, granite3.3 = 0.56, versus claude-opus-4.5 = 0.94 and gemini-2.0-flash = 0.98. Models with $H/\log|\mathcal{D}|<0.6$ achieve determinism primarily through decision concentration rather than task understanding.
Portfolio Constraint Results.
Table 5 shows results across 7 models on the multi-constraint portfolio evaluation task (50 trades).
Table 5: Portfolio Constraint benchmark (50 trades at T=0.0T{=}0.0). Dec.Det and Sig.Det are case-level metrics (fraction of 50 trades with all runs identical).
Model Tier Dec.Det (%) Sig.Det (%) Acc (%) Tools/Run
qwen2.5:7b Tier 1 100.0 100.0 20.0 4.7
gpt-oss:20b Tier 1 62.0 46.0 33.8 3.4
gemini-2.0-flash Frontier 96.0 70.0 62.7 3.6
gemini-2.5-pro Frontier 50.0 21.7 57.2 2.8
claude-sonnet-4 Frontier 86.0 58.0 66.7 4.0
claude-opus-4.5 Frontier 78.0 52.0 40.0 4.5
Text-only baseline (no tool invocation):
granite3.3§ Tier 1 82.2 100.0 51.1 0.0
§granite3.3 does not invoke tools; decisions are text-only (not comparable to tool-calling models).
Portfolio Constraint reveals the starkest determinism-accuracy divergence. qwen2.5:7b defaults to “modify” on 82% of trades (perfect determinism, 20% accuracy), while claude-sonnet-4 distributes decisions across approve/modify/reject with 86% determinism and 66.7% accuracy—the highest in this benchmark. gpt-oss:20b shows the lowest decision determinism of any model (62%), with visible run-to-run decision instability on ambiguous trades.
DataOps Exception Results.
Table 6 shows results from the DataOps Exception benchmark, evaluating data quality exception handling in financial pipelines.
Table 6: DataOps Exception benchmark (50 exceptions at T=0.0T{=}0.0). Decisions: auto_fix, escalate, or quarantine. Dec.Det and Sig.Det are case-level metrics. granite3.3 is excluded (does not invoke tools required for DataOps); gemini-2.5-pro is excluded due to rate limiting during data collection.
Model Tier Dec.Det (%) Sig.Det (%) Acc (%) Tools/Run
qwen2.5:7b Tier 1 94.0 94.0 42.2 4.1
gpt-oss:20b Tier 1 84.0 74.0 43.2 3.4
gemini-2.0-flash Frontier 94.0 70.0 34.7 3.5
claude-sonnet-4 Frontier 84.0 26.0 14.0 5.0
claude-opus-4.5 Frontier 64.0 20.0 24.0 4.9
DataOps Exception is the most challenging benchmark for all models. Even qwen2.5:7b drops from perfect to 94% decision determinism, and its accuracy (42.2%) reflects genuine task difficulty. Claude models show the lowest signature determinism (20–26%), exploring highly diverse tool paths across runs. Notably, claude-sonnet-4’s accuracy (14%) is substantially lower than on other benchmarks—it defaults to “quarantine” on 94% of cases, a conservative failure mode that is deterministic but incorrect. For audit-critical deployments, errors that are at least consistent may be preferred over accurate but unpredictable variance.
Stress Test Results.
Preliminary stress-test results from a v1 pilot (10-case subset, Compliance Triage only) are reported in Appendix E. Because several stress scenario values are projected rather than measured, we do not treat these as part of the core empirical contribution. The pilot data suggest that schema-first architectures with Tier 1 models maintain near-perfect determinism under redeployment, data quality faults, and market shocks, but full 50-case stress experiments are needed to confirm these patterns.
7 Evaluation-Based Deployment Considerations
Based on these findings, we outline tier-specific evaluation considerations for regulated financial environments. Determinism serves as a gating property—necessary for audit readiness but not sufficient evidence of production readiness. These considerations are informed by our benchmark results and should be validated on institution-specific workloads before adoption; benchmark accuracy levels in this study do not by themselves justify deployment.
7.1 Tier 1 Agents (7–20B Parameters)
Evaluation profile: Deterministic under standard conditions.
  • Tools: Any complexity—these models handle multi-tool workflows while maintaining determinism
  • Audit: Sample 5% of trajectories for human review
  • SLO (Service Level Objective): 100% trajectory determinism at T=0.0
  • Validation: Standard sample sizes ($\phi=1.0\times$)
Use Cases:
AML (Anti-Money Laundering)/Compliance triage, regulatory reporting, structured data extraction, any workflow requiring audit replay.
7.2 Tier 2 Agents (40–70B Parameters)
Evaluation profile: Moderate drift requiring guardrails.
  • Tools: Schema-constrained only—avoid free-form tool arguments
  • Audit: Sample 20% of trajectories
  • SLO: 95% trajectory determinism
  • Validation: Increased samples ($\phi=1.8\times$)
Use Cases:
Client communications with human oversight, research assistance, moderate-risk advisory workflows.
7.3 Frontier Models (Claude Opus, Gemini Pro)
Evaluation profile: High variance requiring human oversight.
  • Tools: Full capability, but decisions require human approval
  • Audit: Human review of all decisions before execution
  • SLO: Not applicable (human serves as determinism guarantee)
  • Validation: Per-decision validation
Use Cases:
Research and analysis, complex advisory, capability exploration. These models offer superior task performance but may not reliably satisfy autonomous audit replay requirements.
7.4 Tier 3 Agents (120B+ with MoE/Reasoning)
Evaluation profile: High drift; not suitable for autonomous compliance tasks.
  • Tools: Experimental use only
  • Audit: Full trajectory audit required
  • SLO: Do not deploy for compliance-critical tasks
  • Validation: Not viable ($\phi=3.7\times$ makes validation impractical)
Use Cases:
Capability research, proof-of-concept demonstrations. Not suitable for any production deployment requiring regulatory examination.
Figure 4: Validation sample scaling factor ($\phi$) by model tier. Tier 3 models require $3.7\times$ the validation samples of Tier 1 models to achieve equivalent statistical reliability, making validation economically impractical for compliance-critical deployments.
7.5 Evaluation Matrix
Based on the empirical findings across tiers and architectures, Table 7 provides evaluation-informed considerations for common financial use cases. The guidance prioritizes determinism requirements over raw accuracy, aligning model selection with regulatory constraints and operational risk tolerance.
Table 7: Evaluation-informed model selection considerations by use case.
Use Case Model Tier Architecture Rationale
AML/Compliance Triage Tier 1 (7–20B) Schema-first Audit replay required
Portfolio Constraints Tier 1 + fine-tune Schema-first Structured + domain tuning
Regulatory Reporting Tier 1 Schema-first Zero tolerance for variance
Research/Advisory Frontier Unconstrained Accuracy prioritized, HITL
Client Communications Tier 2 Schema-first Moderate det. + oversight
The matrix reflects a key observation from our benchmarks: audit-critical tasks (AML, compliance, regulatory reporting) favor Tier 1 models with schema-first architectures, accepting lower accuracy in exchange for reproducibility. Conversely, advisory workflows can leverage frontier models where superior accuracy justifies human-in-the-loop oversight to manage non-deterministic variance.
8 Discussion
8.1 Why Determinism Over Accuracy
A natural question arises: why accept 30–40% accuracy (Tables 4, 5) for perfect reproducibility?
In regulated settings, non-deterministic accuracy cannot be validated or audited. An agent producing 75% accuracy with 50% determinism yields statistically unreliable validation: the observed accuracy on any held-out sample may not reflect production behavior due to trajectory variance. Determinism defines the feasible region in which accuracy improvements are meaningful. To be clear: we do not advocate deploying 30% accurate models; we argue that until a model achieves high determinism, accuracy metrics are not trustworthy for compliance validation.
One practical implication suggested by these results is a staged evaluation path: establish audit-ready baselines with deterministic Tier 1 models, then evaluate accuracy improvements through domain-specific fine-tuning—which may preserve determinism while increasing capability—and reserve frontier models for human-in-the-loop advisory workflows where variance is acceptable because a human serves as the consistency guarantee.
8.2 Connection to Industry Standards
The framework aligns with emerging industry practice:
Financial Services Requirements.
As Anthropic’s Head of Financial Services notes in an industry interview: “In financial services, people do not have the luxury of inconsistent outputs” (Robinson, 2026). This practitioner perspective aligns with the regulatory requirements motivating DFAH: evaluation infrastructure to measure whether deployed agents achieve reproducible behavior.
8.3 Regulatory Perspective
From an audit standpoint, replayability and decision reproducibility dominate model optimality. U.S. banking supervisory guidance on model risk management (Board of Governors, 2011), the EU AI Act’s requirements for high-risk system auditability (European Parliament, 2024), and the NIST AI Risk Management Framework’s emphasis on measurable trustworthiness (NIST, 2023) all prioritize consistency under identical inputs over marginal accuracy gains that cannot be explained or reproduced.
Consider a regulatory examination scenario: an examiner requests demonstration that a flagged transaction would be classified identically if the system were re-run. A model achieving 80% accuracy with 50% determinism fails this examination regardless of capability metrics—the examiner cannot be assured the observed decision reflects systematic behavior rather than sampling variance.
The DFAH framework directly addresses this requirement: in our benchmarks, Tier 1 models with schema-first architectures achieved the determinism levels consistent with audit replay requirements.
8.4 Model Size and Determinism
In this evaluation, 7–20B models achieved the highest determinism while maintaining task-sufficient performance. This aligns with emerging findings across modalities: Gan et al. (2025) show that Qwen2.5-7B fine-tuned on motion-specific data can match or exceed larger models on video motion benchmarks.
These results suggest that for tasks prioritizing reproducibility over peak capability, smaller task-optimized models may offer favorable trade-offs. This is not a universal claim however—these findings are specific to the compliance-oriented benchmarks evaluated here. Institutions should validate determinism characteristics on their own workloads before generalizing.
8.5 Limitations
Threat Model Scope.
The framework addresses two failure modes: (1) audit replay failure—inability to reproduce a decision under examination; (2) evidence fabrication—decisions based on hallucinated rather than retrieved facts. Explicitly not addressed: ultimate correctness (a deterministic wrong answer is still wrong), fairness beyond evidence alignment, or adversarial robustness. We make no claims about adversarial settings where inputs are crafted to induce inconsistency.
Statistical Power.
Determinism is evaluated over 3–8 runs per configuration. Rare drift events (occurring <<5% of runs) may be underestimated. With $n=21$ model-benchmark configurations, the study is underpowered to detect moderate correlations (post-hoc power: 7.7% at $\alpha=0.05$); 80% power would require $n\geq 611$ configurations. The null correlation finding (r=0.11r=-0.11) should therefore be interpreted as “no evidence of a relationship” rather than “evidence of no relationship.”
Benchmark Generalizability.
The three benchmarks (compliance triage, portfolio constraints, DataOps exceptions) span structured and semi-structured financial tasks but do not cover open-ended advisory, multi-document reasoning, or cross-lingual workflows. The evaluation uses simulated tool contexts; production environments may introduce additional variance from network latency, API versioning, and data staleness.
Measurement Bias.
Determinism metrics treat any output difference as a failure, even semantically equivalent rephrasings. The faithfulness heuristic is first-order (individual claim-evidence alignment) and may miss relational inversions (Sudjianto, 2026). Tool-call order is treated as significant for signature determinism, though order may not affect correctness.
Ethical Considerations.
Deterministic agents can amplify biases present in retrieved evidence—a consistent wrong answer may be worse than an inconsistent one if it escapes review. High-stakes decisions (e.g., AML triage) warrant human oversight regardless of determinism scores. Our framework should complement, not replace, human judgment in compliance workflows.
8.6 Future Directions
Geometric Approaches to Faithfulness. The evidence-grounding heuristic in this work is a first-order metric: it verifies whether individual claims align with individual evidence items. Recent work on Semantic Plane Alignment (Sudjianto, 2026) demonstrates that second-order metrics—measuring whether relationships between claims match relationships between evidence—can detect relational inversions invisible to first-order approaches. For compliance contexts where causal direction matters (e.g., “rate hike caused inflation” vs. “inflation caused rate hike”), extending the DFAH faithfulness metric with bivector-based alignment is a promising direction.
Beyond geometric faithfulness, natural extensions include multimodal agent evaluation as financial workflows incorporate document images and charts, and domain-specific fine-tuning experiments to determine whether Tier 1 models can improve task accuracy while preserving their determinism characteristics.
9 Conclusion
In this study, we measured reproducibility in tool-using LLM agents deployed in financial services using the Determinism-Faithfulness Assurance Harness (DFAH). Based on 4,705 agentic runs across 7 models and 3 financial benchmarks, we find that determinism and accuracy show no detectable correlation and must be measured independently.
Key Findings.
  1. 1.
    Decision determinism and task accuracy show no detectable correlation (r=0.11r=-0.11, 95% CI [0.49,0.31][-0.49,0.31], $n=21$), indicating that single-metric evaluation is insufficient for deployment assurance
  2. 2.
    Small models (7–20B) achieve near-perfect determinism (94–100%) through rigid pattern matching, at the cost of accuracy (20–42%)
  3. 3.
    Frontier models show moderate determinism (50–96%) with higher but variable accuracy (14–69%), exploring diverse tool paths to reach decisions
  4. 4.
    No model occupies the “high determinism and high accuracy” quadrant, supporting the need for DFAH’s multi-dimensional approach
  5. 5.
    The “same conclusion, different reasoning” pattern—where frontier models converge on decisions while varying their tool paths—is, to our knowledge, a finding not previously documented in agentic evaluation
Practical Considerations.
For compliance-critical deployments, our results suggest that Tier 1 models with schema-first architectures merit evaluation: in our benchmarks, these configurations achieved the determinism levels consistent with audit replay requirements. Frontier models may be better suited for human-in-the-loop advisory workflows where their superior task performance justifies the oversight overhead. Institutions should validate these patterns on their own workloads before deployment.
Industry Alignment.
The framework implements the trial/trajectory/grader evaluation structure emerging in agent evaluation practice (Anthropic, 2026), connecting the metrics to the passk requirements specific to compliance contexts. The three benchmark tasks (150 total test cases) and open-source stress-test harness enable institutions to validate agent deployments against these requirements before production exposure.
Acknowledgments
The author acknowledges the assistance of AI coding harnesses (open-source and frontier LLMs) in developing evaluation code, analyzing experimental data, and preparing this manuscript. All experimental results, analysis, and scientific claims were produced by the human author. All remaining errors are the author’s own. The views expressed herein are those of the author and do not necessarily reflect the views of his employer.
Correspondence: raffi.khatchadourian1@ibm.com
Appendix A Detailed Benchmark Specifications
A.1 Compliance Triage Alert Categories
The 50 compliance test alerts span the following risk categories:
  • Sanctions-related (8 alerts): Direct sanctions hits, sanctions-adjacent entities, high-risk country destinations
  • PEP exposure (4 alerts): Politically exposed persons, government official payments, FCPA risk
  • Structuring patterns (6 alerts): Just-under-threshold transactions, velocity spikes, layered transfers
  • High-value goods (7 alerts): Art, jewelry, luxury vehicles, real estate transactions
  • High-risk sectors (10 alerts): Gaming, weapons, crypto, extractive industries, pharmaceuticals
  • Standard business (15 alerts): Intercompany transfers, vendor payments, payroll (dismiss baseline)
A.2 Portfolio Constraint Violation Types
The 50 portfolio test cases cover:
  • Position limit violations (12 cases): Exceeding 5% single-stock concentration
  • Sector cap violations (6 cases): Technology sector exceeding 25% cap
  • Liquidity violations (8 cases): Small-cap stocks with insufficient trading volume
  • Cash reserve violations (3 cases): Depleting below 2% minimum cash requirement
  • Clean approvals (21 cases): Trades satisfying all constraints
A.3 DataOps Exception Types
The 50 data exception test cases include:
  • Format errors (15 cases): Date formats, currency codes, numeric parsing, ticker symbols
  • Business rule violations (18 cases): Negative prices, invalid ratios, constraint violations
  • Reference data mismatches (10 cases): Ticker mappings, CUSIP/ISIN lookups, LEI validation
  • Missing required fields (7 cases): CUSIPs, settlement dates, accrued interest
Appendix B Stress Test Perturbation Details
The stress-test harness applies four perturbation types:
Redeploy:
Same model redeployed after deployment cycle refresh (e.g., container restart, model reload). Tests intra-model consistency across deployment events.
Temporal Shift:
Use context data from 6 months prior (stale filings, outdated market data). Tests temporal robustness. (Not included in this evaluation; planned for future work.)
DQ Fault:
Inject data quality issues: 10% missing fields, format inconsistencies, null values. Tests error handling.
Market Shock:
Simulate 3-sigma volatility spike in market data. Tests behavior under extreme conditions.
Each perturbation is applied independently, with 8–16 runs per configuration to measure determinism degradation:
$\Delta\text{Det}_{p}=\text{Det}_{\text{baseline}}-\text{Det}_{p}$
A robust agent configuration should maintain $\Delta\text{Det}_{p}<10\%$ across all perturbation types.
Appendix C API Reproducibility Caveats
Seed parameters (seed=42) were passed to all API calls where supported. However, many hosted APIs do not guarantee seed determinism:
  • Local inference (Ollama): Full seed control; reproducible.
  • Anthropic API: Seed parameter not supported; best-effort T=0.
  • Google Gemini API: Seed parameter accepted but not guaranteed.
  • watsonx API: Seed parameter supported; variability observed for some models (granite on watsonx vs Ollama).
This API-level variance explains why identical models (e.g., granite-3-8b) show different determinism rates across providers, and why frontier models (Claude, Gemini) show non-zero drift even at T=0.
Appendix D Statistical Precision of Determinism Measurements
All determinism percentages in the main text are computed from 8–16 runs per configuration. For binomial proportion estimates with $n=8$ runs:
  • 100% determinism (8/8 identical): 95% CI $\approx$ [63%, 100%] (Wilson score)
  • 90% determinism (7/8 identical): 95% CI $\approx$ [56%, 100%]
  • Tier comparisons use $n=24$ aggregate runs (3 tasks $\times$ 8 runs), yielding tighter intervals
  • Agentic benchmarks aggregate 80 runs/model (10 alerts $\times$ 8 runs), yielding CIs such as [82%, 98%] for 90% observed determinism (9/10 alerts)
The experimental design prioritizes detecting instability (any non-zero drift) rather than precise effect sizes. Configurations showing 100% determinism across all runs provide strong evidence of reproducibility, while those below 90% clearly indicate deployment risk.
D.1 Coverage and Power
In v2, agentic experiments use the full 50-case test sets across all benchmarks:
Table 8: Experimental coverage of benchmark tasks (v2). Models vary in runs/case (3 for API models, 8 for local models) due to cost constraints.
Benchmark Cases Models Runs/Case Total Runs
Compliance Triage 50 7 3–8 1,782
Portfolio Constraint 50 7 3–8 1,748
DataOps Exception 50 5 3–8 1,175
Total 4,705
Plus 2 partial-coverage models (deepseek-r1:8b, mistral:7b; 2 cases each) contributing to
Figure 2 but excluded from Table 4 due to insufficient coverage.
granite3.3 and gemini-2.5-pro excluded from DataOps due to tool-calling incompatibility
and rate limiting, respectively.
Full 50-case coverage enables detection of moderate effects across all task types. Two API models (Claude Opus 4.5, Gemini 2.5 Pro) have partial coverage due to credit exhaustion and rate limiting, respectively; their available data is included in analysis.
Appendix E Pilot Stress-Test Results
Table 9 and Figure 5 report pilot stress-test results from v1 (10-case subset, Compliance Triage only). Baseline values are measured; stress scenario values are projected and should be treated as preliminary.
Table 9: Run-level decision determinism (%) under stress scenarios on Compliance Triage benchmark (v1 pilot data, 10-case subset). Baseline values are measured; stress scenario values (Redeploy, DQ Fault, Vol. Shock) are projected based on observed degradation patterns and marked with . Full 50-case agentic results are in Tables 46.
Config Baseline Redeploy DQ Fault Vol. Shock
unconstrained + gpt-oss:20b 5.0 5.0 2.5 5.0
schema-first + gpt-oss:20b 40.0 37.5 32.5 40.0
schema-first + qwen2.5:7b 100 100 93.8 100
schema-first + claude-opus-4.5 100 87.5 100 100
Local (Ollama), API (Anthropic), Projected
Figure 5: Decision determinism under stress conditions on the Compliance Triage benchmark (v1 pilot, 10-case subset). The schema-first architecture with Tier 1 models maintains near-perfect determinism across stress scenarios. The red dashed line indicates the 95% compliance threshold. Stress scenario values marked are projected, not measured.
Appendix F Validation Scaling Factor Derivation
The validation scaling factor $\phi(m)$ referenced in Section 4 quantifies how much larger a validation sample must be for high-drift models to achieve equivalent statistical reliability as low-drift models.
Drift Variance Estimation.
For each model $m$, we compute the drift rate $\delta_{c}$ per configuration $c$ as the fraction of runs producing non-identical outputs:
$\delta_{c}=1-\text{DecDet}_{c}$
The drift variance for model $m$ is then:
$\sigma^{2}_{\text{drift}}(m)=\frac{1}{|C_{m}|}\sum_{c\in C_{m}}(\delta_{c}-\bar{\delta}_{m})^{2}$
where $C_{m}$ is the set of configurations for model $m$ and $\bar{\delta}_{m}$ is the mean drift rate.
Scaling Factor Computation.
Following Ludwig et al. (2024), the required validation sample size scales with drift variance. We define the scaling factor using a fixed reference variance $\sigma^{2}_{\text{ref}}=0.1$ (calibrated to enterprise financial task variance):
$\phi(m)=1+\frac{\sigma^{2}_{\text{drift}}(m)}{\sigma^{2}_{\text{ref}}}$
This ensures $\phi=1$ for deterministic models ($\sigma^{2}=0$) and scales linearly with drift variance for unstable models.
Empirical Values.
From 74 configurations across 12 models:
  • Tier 1: $\sigma^{2}_{\text{drift}}=0.00$ (all 100% determinism), $\phi=1.0\times$
  • Frontier: $\sigma^{2}_{\text{drift}}=0.034$, $\phi=1+0.034/0.1=1.34\times$
  • Tier 2: $\sigma^{2}_{\text{drift}}=0.080$, $\phi=1+0.080/0.1=1.8\times$
  • Tier 3: $\sigma^{2}_{\text{drift}}=0.27$, $\phi=1+0.27/0.1=3.7\times$
The $3.7\times$ factor for Tier 3 models indicates that to achieve the same confidence interval width on a debiased estimate, validation samples must be $3.7\times$ larger than for Tier 1 models. This reflects the measurement error introduced by output instability. Bootstrap 95% CI for $\phi(\text{Tier3})$: $2.9\times,4.8\times]$ based on 1000 resamples of configuration-level drift rates.